1. In an economy, 2 percent of the employed lose their job every month (s = 0.02) and 10 percent of the unemployed find a job every month (f = 0.10).

a. What is the steady-state rate of unemployment?

b. If the economy were at a steady-state unemployment rate and the labour force were 20 million, how many individuals would lose their jobs each month?

Suppose the GDP growth rate is 6 percent and inflation is 2 percent.

If the velocity of money remains constant, what is the change in real money balances?
Answer #1

Q1.

a. Steady state of unemployment rate = 16.67%.

s/(s+f) = 0.02/(0.02+0.1) = 0.1667 = 16.67%

So, the steady state of unemployment rate is 16.67%.

b. Answer= 333,320

Steady state of unemployment = 16.67%

So, at this level, employment rate is 83.33%.

Total workforce is 20,000,000.

Total number of people employed is 83.33%(20,000,000) = 16,666,000

It is given that 2% of the employed loss their job each month. So, at the steady state, the number of people losing job each month would be

0.02(1,666,000) = 333,320

So, 333,320 people lose their job every month when the economy is at steady state.